OpenStudy (anonymous):

The height of an equilateral triangular prism increases by three units. The new volume is more than the original by how much? hree more than the the area of the base three more than the length of the height three times the area of the base three times the length of the height

OpenStudy (campbell_st):

Volume of a prism is base area x height In this question the base area is unchanged. V = A x h is the height increases by 3 V = A(h +3) V = Ah + 3A Hopefullt this helps to find the increase in volume

OpenStudy (anonymous):

three more than the length of the height or three times the length of the height

OpenStudy (campbell_st):

the difference between the two volumes is 3 x the area V=Ah 1st volume V = Ah +3A when the height is increased by 3 units

OpenStudy (anonymous):

wats da answa??

OpenStudy (anonymous):

A.three more than the the area of the base B. three more than the length of the height C.three times the area of the base D. three times the length of the height

OpenStudy (merchandize):

the area of the base remains the same. that appears to be a given. only the height has changed. the volume of a prism is equal to the height * the area of the base. assume the area of the base is y and assume the height is x, then the original volume is equal to x*y. now the height is increased by 3 units. that makes the height equal to (x+3). the new volume is now equal to (x+3)*y which becomes xy + 3y the new volume is xy + 3y the old volume is xy it appears that the new volume is more than the original volume by 3 times the area of the base. i believe that's your answer and it would be worded that way. 3 times the area of the base is the additional volume measured in cubic units. the area of the base is measured in square units.

OpenStudy (campbell_st):

try C

OpenStudy (anonymous):

so wats da answa??

OpenStudy (merchandize):

c

OpenStudy (campbell_st):

please read the offered solutions.... the answer was included in every post